Circular No. 8427

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004fh–2004ft R. J. Foley, University of California at Berkeley; and W. M. Wood-

Vasey, Center for Astrophysics, on behalf of the ESSENCE project (IAUC 8261, 8251, 8238, 8237, 8021, 8034), report the discovery of 13 supernovae on multiple R- and I-band images taken at the Blanco 4-m telescope during Oct. 4–22 by project members. Spectroscopy was obtained on Oct. 10–22 with the Gemini North and South telescopes (+ GMOS) and the Keck I telescope (+ LRIS), showing each to be of type Ia (except for 2004fp, which is of type II). Each new supernova is located between 1′′–3′′ from the center of its host galaxy.

SN  2004 UT  α_{2000}  δ_{2000}  R  z
2004fh Oct. 9 23^{h}28^{m}27.^{s}20 − 8^{o}36′55′′.2 20.9 0.2
2004fi Oct. 9 23 29 45.35 − 8 54 36.3 21.5 0.3
2004fj Oct. 11 1 09 51.07 + 0 27 21.0 22.0 0.2
2004fk Oct. 11 1 13 35.84 − 0 09 27.6 22.9 0.3
2004fl Oct. 11 23 26 57.92 − 9 37 19.1 21.9 0.5
2004fm Oct. 11 23 26 58.14 − 9 37 19.4 22.5 0.4
2004fn Oct. 11 23 30 20.12 − 9 58 30.7 22.5 0.6
2004fo Oct. 13 1 13 28.97 + 0 35 16.3 22.3 0.5
2004fp Oct. 13 2 29 53.01 − 9 01 16.6 21.6 0.2
2004fq Oct. 13 23 27 45.64 − 8 31 12.8 21.5 0.5
2004fr Oct. 15 2 28 43.77 − 8 54 24.0 22.2 0.2
2004fs Oct. 15 2 31 19.95 − 8 49 21.7 22.6 0.5
2004ft Oct. 17 2 33 32.63 − 8 09 34.1 22.4 0.5

COMETS C/2004 T4, C/2004 T5, C/2004 U3 (SOHO) Additional SOHO sungrazers (continuation to IAUC 8426):

Comet  2004 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2004 T4 Oct. 4.904 12^{h}26.^{m}4 − 5^{o}57′ C3/2 HS 2004-V05
C/2004 T5  5.488 12 25.6 − 6 22 C3/2 RK 2004-V05
C/2004 U3  18.580 13 32.1 − 7 55 C2  KB 2004-V05

C/2004 T4 and C/2004 T5 are Kreutz sungrazers, while C/2004 U3 belongs to the Meyer group.

2004 November 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green